OpenStudy (anonymous):

The chemical reaction for making benzene from carbon and hydrogen is: 6C+3H2>C6H6 If 72 grams of carbon and 6 grams of hydrogen react completly, how many benzene will be formed? Explain answer?

OpenStudy (anonymous):

First find moles of each compound Moles= given weight/molecular weight C=72/12=6. H=6/2=3. Now from the equation it is obvious that 6 moles of carbon react with 3 moles of hydrogen to give one mole of benzene. So you have one mole of benzene which weighs 1*78=78 g.